OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x-2/x^2+3x-10

OpenStudy (lgbasallote):

\[\frac{2}{x^2 + 3x + -10} \implies \frac{2}{(x+5)(x-2)}\]

OpenStudy (lgbasallote):

i just factored out x^2 + 3x - 10

OpenStudy (anonymous):

this is how the problem looks and I don't know if that is the answer you came up with.. but it's not one of the answers available... & I am not really understanding this at all.

OpenStudy (anonymous):

@lgbasallote there is x - 2 in the numerator..

OpenStudy (anonymous):

\[\frac{(x-2)}{x^2 + 3x + -10} \implies \frac{(x-2)}{(x+5)(x-2)}\] Cancel what you can cancel and you will have your answer..

OpenStudy (anonymous):

Can you solve it or not @mlb8482 ??

OpenStudy (anonymous):

sorry i was working on another problem just one moment and i will see thank you!

OpenStudy (anonymous):

1/x+5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes you are right dear..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

whoo

OpenStudy (anonymous):

Welcome dear..